OpenStudy (anonymous):

which of the following is a factor of 7x^2 -12x-4? a) x+2 b)none of the above c)7x-2 d)x-4

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

7x^2-12x-4=0 7x^2-14x+2x-4=0 7x(x-2)+2(x-2)=0 (7x+2)(x-2)=0

OpenStudy (anonymous):

ans:7x-2 i guess

OpenStudy (anonymous):

@xen7 After showing that the answer is B by factoring, you then say the answer is c??

OpenStudy (anonymous):

The answer is B

OpenStudy (anonymous):

Yes, we know, I am asking zen why he proves b and answers c.

OpenStudy (anonymous):

Ohhk

OpenStudy (anonymous):

I m sorry the answer is B. Actually I'm sleepy...

OpenStudy (anonymous):

b)none of the above